6.12 The average length of steel nails is 5 centime ters, with a standard deviation of 0.05 centimeters. Assuming that the lengths are normally distributed, what percentage of the nails are




(b) between 4.95 and 5.05 centimeters in length?




(c) shorter than 4.90 centimeters?




(a) longer than 5.05 centimeters?

1
Expert's answer
2022-01-04T04:52:06-0500

Let "X=" the length of steel nail: "N(\mu, \sigma^2),"

Given "\mu=5cm, \sigma=0.05cm."

(a)


"P(X>5.05)=1-P(X\leq 5.05)"

"=1-P(Z\leq \dfrac{5.05-5}{0.05})=1-P(Z\leq 1)"

"\approx 0.1587"

"15.87\%"


(b)


"P(4.95<X<5.05)"

"=P(X<5.05)-P(X\leq 4.95)"

"=P(Z<\dfrac{5.05-5}{0.05})-P(Z\leq \dfrac{4.95-5}{0.05})"

"=P(Z<1)-P(Z\leq -1)"

"\approx0.84134- 0.15866"




"\approx0.6827"

"68.27\%"


(c)


"P(X<4.90)=P(Z<\dfrac{4.90-5}{0.05})=P(Z<-2)"

"\approx 0.02275"

"2.275\%"